# Concept-1K: A Novel Benchmark for Instance Incremental Learning

### v2

@@ Abstract
Incremental learning (IL) is essential to realize the human-level intelligence in the neural network. However, existing IL scenarios and datasets are unqualified for assessing forgetting in PLMs, giving an illusion that PLMs do not suffer from catastrophic forgetting. To this end, we propose a challenging IL scenario called instance-incremental learning (IIL) and a novel dataset called Concept-1K, which supports an order of magnitude larger IL steps. Based on the experiments on Concept-1K, we reveal that billion-parameter PLMs still suffer from catastrophic forgetting, and the forgetting is affected by both model scale, pretraining, and buffer size. Furthermore, existing IL methods and a popular finetuning technique, LoRA, fail to achieve satisfactory performance. Our study provides a novel scenario for future studies to explore the catastrophic forgetting of PLMs and encourage more powerful techniques to be designed for alleviating the forgetting in PLMs. The data, code and scripts are publicly available  [^1] https://github.com/zzz47zzz/codebase-for-incremental-learning-with-llm.
Concept-1K: A Novel Benchmark for Instance Incremental Learning
@@ 1 Introduction
Incremental Learning (IL) aims to equip models with new knowledge without forgetting the old knowledge. It is of critical importance for research on the verge of human-level intelligence. Recent years have witnessed significant breakthroughs of Pretrained Language Models OpenAI (2023) (PLMs) in not only downstream tasks in natural language processing but also in real-world applications. Although there is a wealth of research on incremental learning with PLMs, we still lack a fundamental understanding of the incremental learning ability of PLMs. For instance, we can not clearly answer the questions “Do PLMs still suffer from catastrophic forgetting?”, “Do larger PLMs forget less than the smaller ones?”, “What is the role of pretraining in incremental learning?”. The root cause is the lack of suitable IL benchmarks designed for PLMs.
Figure 1: The illustration of Instance Incremental Learning (IIL).
There are primarily three IL scenarios for PLMs: class-incremental learning (CIL), task-incremental learning (TIL), and continual pretraining (CPT). However, none of them are appropriate to evaluate the IL ability of PLMs. CIL is designed for classification tasks such as text classification, and its goal is to learn new classes incrementally. On the one hand, state-of-the-art (SOTA) PLMs with billion-level parameters are overqualified for the above classification tasks with only dozens of categories. On the other hand, the pretraining corpus is likely to contain the knowledge required for the classification tasks (data leakage issue). As shown empirically by Zheng et al. (2023), sequential training frozen PLMs with expanding classifiers yields comparable or even superior performance with SOTA IL methods. TIL aims to learn new tasks incrementally Sun et al. (2020); Qin and Joty (2022). Apart from the data leakage issue, the diversity of tasks and orders across research makes it difficult to readily and fairly compare IL algorithms. The last scenario CPT aims at continual pretraining models on the corpus from different domains. However, the evaluation relies on the performance of downstream tasks, where we can hardly identify what knowledge is learned or forgotten.
In this paper, we propose a novel incremental learning scenario called Instance Incremental Learning (IIL), for benchmarking the IL ability of SOTA PLMs. Unlike the IL scenario mentioned above, IIL regards each concept as an instance and is more practical and challenging for existing PLMs. Specifically, we are motivated by the human learning process and expect PLMs to learn new concepts incrementally without forgetting. For example, in Figure 1, humans can learn new concepts that are constantly emerging, such as “Metaverse” and “Quantum Computing”. After learning more concepts such as “Web3.0” and “Non-Fungible Token”, humans will not immediately forget the previously learned concepts such as “Metaverse”. In contrast, we will later empirically show that PLMs with billions of parameters fail to recall the knowledge of old concepts.
Additionally, we present a dataset called Concept-1K for the IIL scenario. Concretely, we utilize the vast knowledge of GPT4 OpenAI (2023) to generate recently emerged concepts from six domains, including economy, culture, science and technology, environment, education, and health and medical, resulting in a total of 1023 concepts. Next, we prompted GPT4 to construct triplets for each concept with the relation in ConceptNet Speer et al. (2017) where each triplet instantiates a knowledge point. Finally, we obtain 16653 training-test QA pairs corresponding to 16653 knowledge points from 1023 novel concepts.
We conduct experiments on different scales of PLMs with various pretraining steps, buffer size, and finetuning methods under the IIL scenario. The results show that PLMs suffer from catastrophic forgetting seriously and have a huge gap to human beings. In addition, we reveal that existing rehearsal-free IL methods fail to achieve superior performance. Furthermore, we find that in-context learning fails to incrementally adapt PLMs to novel concepts. In summary, the contribution is two-fold:
  • We propose a novel incremental learning scenario called IIL, and a dataset called Concept-1K for better evaluating the IL ability of PLMs.
  • Based the experiments on Concept-1K, we reveal that PLMs still suffer from catastrophic forgetting and existing techniques fail to achieve superior performance when no old data is stored.
2 Instance Incremental Learning
@@
Problem Formulation. Formally, IIL aims to learn a model $f_{\theta}:\mathbf{x}\rightarrow\mathbf{y}$ from the sequence of concepts $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n},\cdots,\mathcal{C}_{N}\}$, where $N$ is the number of concepts, and both input $\mathbf{x}$ and output $\mathbf{y}$ are natural language. The $n$-th concept $\mathcal{C}_{n}$ contains $M_{n}$ training-test pairs $\mathcal{D}^{(n)}=\{\mathbf{x}_{i}^{(n),train},\mathbf{x}_{i}^{(n),test},\mathbf{y}_{i}^{(n)}\}_{i=1}^{M_{n}}$, where $\mathbf{x}_{i}^{(n),train}$ and $\mathbf{x}_{i}^{(n),test}$ are the training and test inputs, and $\mathbf{y}_{i}^{(n)}$ is the target output of them. Each training-test pair corresponds to the same knowledge point about the concept $\mathcal{C}_{n}$. For practical training and evaluation, we evenly divide $N$ concepts into $T$ ($T\leq N$) groups corresponding to $T$ incremental steps. The model will be evaluated after learning each group of concepts.
@@
Metric. We adopt four evaluation metrics for IIL, including Memorization Accuracy (MA), Memorization Forgetting rate (MF), Generalization Accuracy (GA), and Generalization Forgetting rate (GF). Specifically, MA and MF measure how much knowledge of training samples is memorized and forgotten. GA and GF measure how much knowledge is generalized to the test samples and is forgotten. The formal definition is provided in Appendix A.
@@
Table 1: Comparison between Concept-1K and the widely-used datasets for incremental learning with PLMs. Concept-1k supports an order of magnitude larger incremental learning steps than existing ones.
@@
Concept-1K 1023 Question and Answering

### v3

# Can LLMs Learn New Concepts Incrementally without Forgetting?

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success across various tasks, yet their ability to learn incrementally without forgetting remains underexplored. Incremental learning (IL) is crucial as it enables models to acquire new knowledge while retaining previously learned information, akin to human learning. Existing benchmarks for IL are insufficient due to data leakage issues and the overqualification of LLMs. To address these challenges, we introduce Concept-1K, a novel dataset comprising 1,023 recently emerged concepts across diverse domains. The concepts in Concept-1K are discrete, interpretable units of knowledge that allow for fine-grained analysis of learning and forgetting processes. Using Concept-1K as a testbed, we aim to answer the question: “Can LLMs learn new concepts incrementally without forgetting like humans?” Our investigation reveals that LLMs still suffer from catastrophic forgetting and that LoRA, despite fine-tuning fewer parameters, may lead to more forgetting on training data. Additionally, we explore the roles of in-context learning, model scale, buffer size, and pretraining in IL performance. These findings highlight the strengths and limitations of LLMs in IL scenarios and provide a robust benchmark for future research. The data, code and scripts are publicly available  [^1] https://github.com/zzz47zzz/codebase-for-incremental-learning-with-llm.
Can LLMs Learn New Concepts Incrementally without Forgetting?
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved remarkable success, exhibiting human-level performance on various professional and academic benchmarks OpenAI (2023). Numerous studies have investigated various abilities of LLMs, such as reasoning Wei et al. (2022), programming Chen et al. (2021), and planning Yao et al. (2022). However, a crucial human ability, incremental learning (IL) (also known as continual learning), remains less explored in LLMs.
Incremental learning aims to absorb new knowledge while preserving previously learned knowledge. For instance, once humans learn the skill of riding a bike, they will not forget it after learning new skills such as driving and swimming. Naturally, one might wonder, “Since LLMs are so powerful, do they still suffer from forgetting when learning incrementally?”
To answer this question, we first need to find a proper benchmark for evaluating the IL ability of LLMs. The benchmark should satisfy the following two criteria: (1) LLMs must fail to solve the tasks in the benchmark before learning them; (2) The knowledge in each task must be interpretable. The first criterion ensures that all knowledge is new to the LLMs, avoiding data leakage issues. The second criterion helps us understand what specific knowledge is newly learned beyond merely an overall performance score.
Figure 1: The illustration of the proposed Concept-1K. LLMs suffer from catastrophic forgetting when learning new concepts while humans do not.
Table 1: The data leakage issue in popular datasets for IL. The linear probing performance Zheng et al. (2023b) on Topic3Datasets, CLINC150, FewRel, OntoNotes5, and I2B2 before IL training is reported, as well as the test accuracy of Concept-1K before IL training. “/” represents not applicable.
Topic3Datasets CLINC150 FewRel OntoNotes5 I2B2 Concept-1K
Pythia-410m 87.45±0.36 91.05±0.65 74.16±0.11 / / 0.68±0.17
bert-base-cased 88.02±0.56 80.39±0.27 52.18±0.05 52.93±0.48 58.29±0.73 /
However, none of the existing benchmarks satisfy these two criteria simultaneously. Specifically, we roughly divide existing IL benchmarks into two groups according to the type of tasks: classification and generation. Classification benchmarks are widely used in IL studies from the pre-LLM era, including text classification Zhang et al. (2015), named entity recognition Ding et al. (2021), and relation extraction Han et al. (2018). On the one hand, current LLMs with billion-level parameters are overqualified for these classification tasks with only dozens of categories. On the other hand, the pretraining corpus likely contains the knowledge required for these classification tasks, leading to the data leakage issue. As shown empirically by Zheng et al. (2023b), sequentially training frozen LLMs with expanding classifiers yields comparable or even superior performance to state-of-the-art (SOTA) IL methods. Generation benchmarks Zhang et al. (2023c); Wang et al. (2022a) include various tasks such as question generation, style transfer, and wrong candidate generation. However, the data leakage issue remains. In the experiments of Zhang et al. (2023c), training T5 Raffel et al. (2020) on 19 various tasks jointly achieves 42.1% average performance (i.e., upper bound performance), while sequential finetuning achieves 35.7% (i.e., lower bound performance). Further discussion on data leakage issues is provided in Appendix B.
To address these challenges, we construct a dataset called Concept-1K, which satisfies the two criteria for investigating the IL ability of open-sourced LLMs such as LLaMa Touvron et al. (2023). Specifically, Concept-1K minimizes the data leakage issue by selecting recently emerged concepts such as “Metaverse” and “Quantum Computing” from various vertical domains that require domain-specific knowledge to answer. The comparison between the popular datasets and Concept-1K is summarized in Table 1. Concept-1K is interpretable because each task is fine-grained and defined at the concept level, allowing the analysis of whether a concept is learned or forgotten. Additionally, Concept-1K contains 1,023 concepts, supporting an order of magnitude larger incremental learning steps than existing benchmarks, which can push LLMs’ IL ability to their limits.
Using the constructed Concept-1K as a testbed, we aim to answer the question: “Can LLMs learn new concepts incrementally without forgetting, like humans?” The choice of “concept” as the fundamental unit in Concept-1K is deliberate. As shown in Figure 1, concepts are discrete, interpretable units of knowledge that allow for fine-grained analysis of learning and forgetting processes. By focusing on concepts, we can precisely identify what knowledge is acquired, retained, or forgotten, providing clearer insights into the incremental learning abilities of LLMs. Our investigation also delves into how in-context learning, parameter-efficient methods like LoRA Hu et al. (2021), and factors such as model scale, buffer size, and pretraining influence IL performance.
Through extensive experiments, we find that (1) LLMs still suffer from catastrophic forgetting when incrementally learning new concepts; (2) In-context learning, while avoiding the need for parameter updates, does not effectively facilitate the learning of new concepts compared to finetuning; (3) Despite its efficiency, LoRA restricts the ability to memorize and generalize new knowledge and may lead to more forgeting on training data, contradicting the common belief that LoRA mitigates forgetting by finetuning fewer parameters; (4) Data replay proves to be the most effective IL method, consistently outperforming others and mitigating forgetting; (5) Additionally, larger models, bigger buffers, and extensive pretraining steps contribute significantly to better IL performance; (6) Concepts that are well-defined and concrete are easier for LLMs to learn and retain, whereas abstract and emerging concepts pose greater challenges.
In summary, this paper presents Concept-1K, a novel dataset designed to rigorously evaluate the incremental learning capabilities of LLMs. Our findings provide valuable insights into the strengths and limitations of current LLMs in IL scenarios and offer a robust benchmark for future research in this area.
2 Concept-1K
Table 2: Examples of Concept-1K. Each triplet corresponds to a training instance and a test instance. More examples are provided in Table 21.
Domain Concept Triplet Training and Test Input Target Output
What kind of process is Groundwater Recharge?
What are essential for the process of Groundwater Recharge?
What does Groundwater Recharge lead to regarding water tables?
Why is Groundwater Recharge important?
What is the primary factor leading to Sea Level Rise?
What group of professionals study Sea Level Rise?
What is a significant impact of Sea Level Rise on natural habitats?
What instrument is used to measure Sea Level Rise?
What strategy addresses Sea Level Rise?
2.1 Problem Formulation
@@
We consider an incremental scenario where LLMs explicitly learn the knowledge of each concept. Specifically, we aim to train a model $f_{\theta}:\mathbf{x}\rightarrow\mathbf{y}$ from a sequence of concepts $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n},\cdots,\mathcal{C}_{N}\}$, where $N$ is the number of concepts, and both the input $\mathbf{x}$ and output $\mathbf{y}$ are natural language. The $n$-th concept $\mathcal{C}_{n}$ contains $M_{n}$ training-test pairs $\mathcal{D}^{(n)}=\{\mathbf{x}_{i}^{(n),train},\mathbf{x}_{i}^{(n),test},\mathbf{y}_{i}^{(n)}\}_{i=1}^{M_{n}}$, where $\mathbf{x}_{i}^{(n),train}$ and $\mathbf{x}_{i}^{(n),test}$ are the training and test inputs, and $\mathbf{y}_{i}^{(n)}$ is the target output. Each training-test pair corresponds to the same knowledge point about the concept $\mathcal{C}_{n}$.
For instance, in Table 2, the target output for both questions, “What is Groundwater Recharge classified as?” and “What kind of process is Groundwater Recharge?” is “hydrological process”. We expect LLMs to learn the knowledge point “Groundwater Recharge, IsA, HydrologicalProcess” from the training sample and generalize it to answer the rephrased test question correctly. For practical training and evaluation, we evenly divide $N$ concepts into $T$ ($T\leq N$) tasks. The model is evaluated after learning the concepts in each task.
2.2 Evaluation Metric
@@
We adopt four evaluation metrics for Concept-1K: Memorization Accuracy (MA), Memorization Forgetting rate (MF), Generalization Accuracy (GA), and Generalization Forgetting rate (GF). Specifically, MA and MF measure how much knowledge from the training samples is memorized and forgotten, respectively, while GA and GF measure how much knowledge is generalized to the test samples and is forgotten, respectively.
Memorization accuracy is defined as:
$MA=\frac{1}{T}\sum_{t=1}^{T}\mathcal{A}_{t},\quad\mathcal{A}_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i},$ (1)
where $T$ is the number of tasks. $\mathcal{A}_{t}$ represents the average accuracy on the training instances from all learned concepts. $a_{t,i}$ represents the accuracy evaluated on the $i$-th task after training the model incrementally from concepts belonging to task $1,\cdots,t$. The accuracy is calculated as the exact match between the model output and the target output.
Memorization forgetting is computed as the average accuracy on all training instances of all learned concepts:
$MF=\frac{1}{T-1}\sum_{i=1}^{T-1}\left[\max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}\right],$ (2)
where $\max_{j<T}(\{a_{j,i}\}_{j})$ represents the highest accuracy of task $i$ since it has been learned, and $a_{T,i}$ represents the accuracy of task $i$ at step $T$. $\left[\max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}\right]$ computes the decrease in the accuracy of task $i$ when learning the $T$-th task.
Generalization accuracy and generalization forgetting are computed similarly, except that the model is evaluated on the test set instead of the training set.
2.3 Dataset Construction
To avoid data leakage, we collect novel concepts from six domains: economy, culture, science and technology, environment, education, and health and medical. Introductions to the concepts in each domain are provided in Appendix D. Initially, we generate 600 concepts for each domain using GPT-4. We then manually filter out outdated, vague, or imaginary concepts and select the latest, most specific, and most informative concepts, resulting in a total of 1,023 concepts. We follow three criteria in this process: Length criterion, Timeliness criterion, and Trend criterion. Detailed description is provided in Appendix E.
The concept list is provided in Table 22. Next, we use triplets to represent “knowledge” and prompt GPT-4 to construct 20 triplets for each concept with the relations in ConceptNet Speer et al. (2017). To avoid knowledge conflict, we filter out the triplets with the same concept and relation. Additionally, we filter out triplets with relations such as “RelatedTo” and “HasContext” to ensure specificity. Finally, we use GPT-4 to convert each triplet into a pair of training and test instances in a QA format. Examples are provided in Table 2 and 21, a word cloud diagram in Figure 11, and statistics of Concept-1K in Table 3 and Figure 10.
@@ 2.3 Dataset Construction
Table 3: Comparison between Concept-1K and widely-used datasets for incremental learning with LLMs. Concept-1K supports an order of magnitude larger incremental learning steps than existing ones.
@@
Concept-1K 1023 Question Answering
2.4 Comparison with Existing Datasets
2.4.1 Concept-1K Minimizes Data Leakage
Concept-1K is designed to minimize data leakage by focusing on novel concepts that emerged after January 2022. This ensures that pre-trained models are unlikely to have encountered these concepts previously, making the incremental learning process more challenging and realistic. The zero-shot performance of models such as GPT-4, GPT-3.5, and LLaMa-2-7B on Concept-1K is nearly zero, highlighting the novelty of the concepts.
2.4.2 Concept-1K Defined as Instance-Level Incremental Learning
Unlike other datasets, which are often designed for task-level incremental learning, Concept-1K is constructed under a new scenario called Instance-level Incremental Learning (IIL). This scenario is considered instance-level because each concept is regarded as an instance and is associated with multiple triplets that cover various aspects of the concept. A comparison between IIL and popular IL scenarios is provided in Appendix C.
2.4.3 Concept-1K Supports More Incremental Tasks
Compared to existing datasets, Concept-1K supports a significantly larger number of incremental learning steps. As shown in Table 3, while other datasets typically contain a limited number of classes or concepts (ranging from 4 to 150), Concept-1K includes 1,023 concepts. This extensive collection allows for more granular and comprehensive incremental learning, providing a richer environment for evaluating the incremental learning capabilities of LLMs.
Table 4: Semantic Diversity in Concept-1K
Concept Name Question and Answer
Intra-domain 0.648 0.758
Inter-domain 0.607 0.704
All 0.613 0.713
Additionally, the concepts, questions, and answers are diverse. We computed the cosine similarity of the average last hidden states of bert-base-uncased Devlin et al. (2019), as shown in Table 4. Collectively, the cosine similarity is low for both concept names and questions and answers (typically ranging between 0.5 and 1.0).
3 Experiments
We split the 1023 concepts in Concept-1K into 10 tasks for incremental learning. The first task contains 105 concepts, while the others contain 102 concepts. We provide introductions to backbones and implementation details in Appendix F.
(a) Memorization Accuracy
(b) Generalization Accuracy
Figure 2: The step-wise performance on Concept-1K. The backbone model is LLaMa-2-7B.
3.1 RQ1: Can LLMs learn new concepts incrementally without forgetting?
Main Findings 1: LLMs still suffer from catastrophic forgetting when incrementally learning new concepts.
We sequentially fully fine-tuned LLaMa-2-7B on 10 tasks from Concept-1K. Before training, we evaluate the LLM on Concept-1K and find that the accuracy on both the training and test data is nearly zero. This indicates that the LLMs lack the knowledge to answer the questions in Concept-1K, thus avoiding the data leakage issue.
Figure 2 shows a clear tendency for the LLMs to forget old concepts’ knowledge when learning new concepts. Specifically, although LLMs achieve 100% memorization accuracy on each new task, the memorized knowledge is gradually forgotten as more tasks are learned. Similarly, the generalized knowledge also diminishes as new knowledge is acquired. Therefore, despite their power, we conclude that LLMs still suffer from catastrophic forgetting when fully fine-tuning on new data.
Table 5: The accuracy of in-context learning on the full Concept-1K dataset. “Rand.”, “Same Conc.”, and “Same Know.” represent that the demonstration samples are selected randomly, or from the instances related to the same concept, or from the instance related to the same knowledge (i.e., the same training-test pair).
1 Shot 5 Shot
Rand. Same Conc. Same Know. Rand. Same Conc. Same Know.
GPT 3.5 6.60 8.20 51.60 8.60 13.00 74.80
GPT 4 10.20 10.40 76.60 7.40 21.80 86.20
3.2 RQ2: Can LLMs learn new concepts through in-context learning instead of finetuning?
Main Findings 2: LLMs hardly learn new knowledge through in-context learning compared to finetuning.
Given the finding that LLMs tend to forget when learning new concepts, we explore in-context learning as a straightforward method that requires no finetuning and does not cause forgetting. For example, Zheng et al. (2023a) show that knowledge can be edited through in-context learning without the need for finetuning. Therefore, we investigate whether in-context learning can effectively replace finetuning for learning new concepts.
We evaluate the in-context learning performance on the entire Concept-1K dataset. Detailed settings and input prompt are provided in Appendix F. Table 5 shows that GPT-4 achieves 86.20% under the “5-shot” and “Same Knowledge” settings, indicating that the training and test instances of Concept-1K share the same knowledge points. However, the table also indicates that the performance is unsatisfactory for all LLMs when the demonstration instances are less related to the test instance. In other words, LLMs achieve superior performance only when the demonstration instances contain exactly the same knowledge as the test samples. Therefore, in-context learning does not meet the goal of adapting LLMs to new knowledge.
Table 5 also shows that the smallest LLM (Pythia-70M) achieves high accuracy under the “1-shot” and “Same Knowledge” settings because small LLMs simply copy the output in the demonstration instance as the final output. Under the “5-shot” and “Same Knowledge” settings, the accuracy of Pythia-70M drops to only 15.02%.
Table 6: The performance of full finetuning (FULL) and LoRA on various backbones.
MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$)
Pythia-410M (FULL) 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16
Pythia-2.8B (FULL) 51.91±0.51 23.18±0.14 42.65±0.64 20.59±0.20
Vicuna 7B (FULL) 77.85±0.71 33.92±0.52 36.35±0.57 22.29±0.49
LLaMa 7B (FULL) 74.69±0.38 30.63±0.46 37.04±0.81 21.05±0.27
Pythia-410M (LoRA) 15.72±0.72 6.58±0.15 30.97±1.15 2.96±0.14
Pythia-2.8B (LoRA) 36.56±0.93 10.93±0.17 83.94±0.64 6.04±0.23
Vicuna 7B (LoRA) 42.28±0.25 16.74±0.33 78.32±0.54 7.69±0.47
LLaMa 7B (LoRA) 41.76±0.27 16.20±0.18 80.55±0.39 8.08±0.31
LLaMa 13B (LoRA) 48.90±0.68 22.67±0.30 74.55±0.46 12.26±0.18
(a) Training Set
(b) Test Set
Figure 3: Comparison of the performance between full finetuning and LoRA on (a) the training set and (b) the test set. The height represents relative performance.
3.3 RQ3: Is LoRA a better choice than full finetuning for IL with LLMs?
Main Findings 3: LoRA is worse than full finetuning and may also lead to more forgetting on training data.
Given the limitations of in-context learning, we turn our attention to LoRA, a method that fine-tunes only a small proportion of parameters. Recently, LoRA has been widely used for designing IL methods or as an experimental setting Zheng et al. (2024). Additionally, Biderman et al. (2024a) argue that LoRA learns less and forgets less.
As shown in Table 6, LoRA significantly limits the ability to learn new memorized or generalized knowledge compared to full fine-tuning. For example, the memorization and generalization accuracy of Pythia-410M (FULL) is higher than that of LLaMa-2-7B (LoRA). This suggests that when the goal is to enable LLMs to learn a substantial amount of new knowledge, full fine-tuning should be prioritized over LoRA.
Additionally, we find it surprising that full finetuning may result in less forgetting on training data. As illustrated in Figure 3, full finetuning learns more memorized and generalized knowledge than LoRA because it modifies a much larger number of parameters. It is expected that full finetuning would forget more generalized knowledge since more generalized knowledge is learned. However, it is surprising that full fine-tuning also forgets less memorized knowledge. This implies that LLMs are more resilient to forgetting when using full finetuning, highlighting the importance of investigating IL in the full finetuning settings instead of LoRA, which is widely adopted in recent IL studies Yang et al. (2024); Ren et al. (2024).
Method MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$) Runtime (min)
SEQ 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16 27
EWC Kirkpatrick et al. (2017) 59.83±0.62 18.09±0.28 62.46±0.69 15.07±0.33 33
LAMOL_g Sun et al. (2020) 58.76±0.53 15.35±0.46 64.64±0.59 13.61±0.47 48
LAMOL_t Sun et al. (2020) 58.29±2.17 15.24±0.37 66.66±2.42 14.05±0.38 48
L2KD Chuang et al. (2020) 28.34±0.29 10.87±0.01 32.45±0.63 8.55±0.42 91
PCLL Zhao et al. (2022) 61.94±1.41 20.06±0.42 63.04±1.57 16.27±0.37 252
LFPT5 Qin and Joty (2022) 0.63±0.04 0.84±0.01 0.04±0.03 0.03±0.01 44
LAMOL_KD Zheng et al. (2023b) 72.33±0.45 18.20±0.37 49.25±0.32 10.61±0.42 59
REPLAY (buffer size=2000) 77.31±0.22 22.48±0.29 46.97±1.50 10.57±0.24 44
REPLAY (buffer size=Alll) 99.01±0.11 25.70±0.44 0.70±0.16 1.44±0.88 110
3.4 RQ4: What is the most effective and efficient method for IL of LLMs?
Main Findings 4: Data replay remains the most effective and efficient method for IL of LLMs.
Given that full finetuning and LoRA both have their own limitations, we explore what the most effective and efficient method for incremental learning of LLMs might be. Data replay is a straightforward approach to IL that stores a small number of samples from previous tasks and optimizes them jointly with new data when learning new tasks. Although numerous IL methods Zheng et al. (2024) have been designed to function without data replay, we find that none of these methods achieve satisfactory performance in our settings.
We compare data replay (REPLAY) with seven SOTA rehearsal-free methods. The introduction of each method is provided in Appendix G. The backbone model used is Pythia-410M. Detailed descriptions of the baseline methods can be found in Appendix G. Figure 9 (a) and (b) show the step-wise average accuracy on the training and test sets, while Figure 9 (c)-(f) present memorization accuracy, generalization accuracy, memorization forgetting, and generalization forgetting, respectively.
Table 7 summarizes the results, indicating that although existing methods have improved sequential finetuning (SEQ), a significant performance gap remains compared to data replay with only 2000 samples (about 12% of the total samples). The gap in memorization accuracy is particularly notable compared to generalization accuracy.
Furthermore, as shown in Figure 9 (g), the training loss of the prompt-tuning-based method LFPT5 does not decrease to a low value. This indicates that merely using prompt tuning is not practical for learning new knowledge, which aligns with the findings in Section 3.3. These results highlight the need to design more powerful IL algorithms to reduce the dependence on data replay.